OpenStudy (anonymous):

The distance covered by an object is governed by the expression x^3+3x^2+2x. If the time taken to cover this distance is given by x + 1, what is the expression that governs the speed of the object?

OpenStudy (anonymous):

(A). x^2+2x (B). x^2+2x-1/x+1 (C). x+2 (D). x^2+1/x+1

OpenStudy (anonymous):

@mathmath333 would you mind.

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{speed =\dfrac{distance}{time} \hspace{.33em}\\~\\ speed =\dfrac{ x^3+3x^2+2x}{ x + 1} \hspace{.33em}\\~\\ ~~~~~~~~~~=\dfrac{x( x^2+3x+2)}{ x + 1} \hspace{.33em}\\~\\ }\end{align}\) factor \(\large \begin{align} \color{black}{(x^2+3x+2) \hspace{.33em}\\~\\ }\end{align}\) and simplify

OpenStudy (anonymous):

I got (x+2)(x+1).

OpenStudy (mathmath333):

yes so substitue that for \(x^2+3x+2\)

OpenStudy (mathmath333):

\(x+1\) will get cancelled

OpenStudy (anonymous):



OpenStudy (anonymous):

so A?

OpenStudy (mathmath333):

\(\Huge \checkmark\)

OpenStudy (anonymous):

Thank you much man your a lifesavor.